OpenStudy (anonymous):

I have 25 % fewer cigarettes than yesterday. Now I have 12. How many did I have yesterday?

OpenStudy (anonymous):

12 + 12*(25%) =12 + 12/4 = 15

OpenStudy (anonymous):

16. 16*.75=12

OpenStudy (anonymous):

Kaiyne is right. Do u want complete solution??

OpenStudy (anonymous):

yea, 16, 25percent of 16 isn't 12.

OpenStudy (anonymous):

*of 15